# Safe MPC Alignment with Human Directional Feedback

### v1

@@ VI-B1 Environment Setup
Figure 11: Drone Navigation Game. (a) The tracking view seen by the user in the drone user study. (b) The fixed view seen by the user in the drone user study. Users can precept the motion of the drone in 3D space with the help of these two views. The $xyz$-axis (red, green, and blue lines), the gate (yellow bars for the edge, orange bars for the outer frame), the drone (white robot model), the local plan outputted from MPC (red curve) and current motion target (yellow balls) are visible to users. in runtime, the user will see one target at a time.
where $\beta_{t}$ is a weighting factor to control or balance the immediate and future constraint satisfaction of robot MPC control. If one emphasizes instant safety, more weights will be given to the initial time step. In our implementation, inspired by the practice in RL, we choose an exponentially decaying $\beta_{t}$ starting at $t=5$, i.e. $\beta_{t}=0,\ t<5$ and $\beta_{t}=0.9^{t-5},\ t\geq 5$ to achieve better predictive safety performance. In (33), each Guassian RBF is defined as
Here, the variables $r_{xy,t}=[r_{x,t},r_{y,t}]^{\mkern-1.5mu\mathsf{T}}$ and $r_{xz,t}=[r_{x,t},r_{z,t}]^{\mkern-1.5mu\mathsf{T}}$ with $r_{x},r_{y},r_{z}$ from $\boldsymbol{x}_{t}$. The $i$th center $c_{xy,i}=[c_{x,i},c_{y,i}]^{\mkern-1.5mu\mathsf{T}}$ or $c_{xz,i}=[c_{x,i},c_{z,i}]^{\mkern-1.5mu\mathsf{T}}$ is defined with $c_{x,i}={x_{\text{gate}}}=9.8$, $c_{y,i}$ and $c_{z,i}$ are evenly chosen on interval $[0,10]$. We set $\epsilon=0.45$. The initial hypothesis space $\Theta_{0}$ for possible $\boldsymbol{\theta}$ in (32) is set as $\boldsymbol{c}_{l}=[-80,\dots,-80]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{h}=[100,\dots,100]^{\mkern-1.5mu\mathsf{T}}$. $\gamma$ is 50. This initialization blocks the drone from advancing to the gate at the beginning of the game, leading to conservative behavior in the learning process. The drone will proceed toward the gate after few corrections.
VI-B2 Human Correction Interface
Human users apply corrections to the drone via a keyboard during the drone flying. The user is informed of the mappings between the key press and the corresponding correction in the game, listed in Table I. The safety constraint (32) in the drone’s MPC policy is immediately updated once a user correction is detected. When the user presses "ENTER", the drone is reset to the initial position and the target will also change.
TABLE I: keyboard commands in uav user study
Key Correction Interpretation
"enter" N/A Reset robot position and switch target.
VI-C Franka Constrained Reaching Game
VI-C1 Environment Setup
As shown in Fig. 12, the goal of the Franka constrained reaching game is to teach a simulated Franka robot arm such that its gripper hand successfully reaches different target positions (shown in cyan ball) through a narrow slot (shown by two yellow bars) with unknown position or geometry. Specifically, the arm starts at a configuration in which the end-effector position $\boldsymbol{r}_{\text{init}}=[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ and altitude $\boldsymbol{q}_{\text{init}}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$. The reaching target position $\boldsymbol{r}^{*}$ will be dynamically chosen from three options $[-0.6,-0.5,0.4]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.6]^{\mkern-1.5mu\mathsf{T}}$. The reaching target attitude is $\boldsymbol{q}^{*}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$. The slot size and position are unknown except its rough x-position which is $x_{\text{slot}}\approx-0.15$.
@@ VI-C1 Environment Setup
Figure 12: Franka Constrained Reaching Game. (a) The starting configuration of the arm in the user study. RGB arrows denote the world $xyz$-frame. (b) Illustration of one successful reach by adjusting the position and altitude of the gripper, long green arrow denotes the motion direction. Users can precept the motion of the robot arm in 3D space by adjusting the view. The bars, the robot arm, and current motion target (cyan ball) are all visible to users.
respectively.
The above end-effector dynamics is discretized with $\triangle t=0.1\mathrm{s}$. In (1), the control horizon is $T=20$. Let $\boldsymbol{x}^{*}=[\boldsymbol{r}^{*}\ \boldsymbol{q}^{*}]$ denote the target end-effector pose. The step cost $c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})$ and terminal cost $h(\boldsymbol{x}_{T})$ in (1) is set to
Similar to Section VI-B, to constrain $r_{z}$ and $\boldsymbol{q}$ for proper height and tilt in order to pass through the slot. We use a set of $N=20$ Gaussian radial basis function (RBFs) to parameter the safety constraint:
with
which is used to constrain the $\mathrm{RBF}_{i}$ only taking effect around $x$ range $[x_{\text{slot}}-\frac{w}{2},x_{\text{slot}}+\frac{w}{2}]$ (recall slot’s $x$ location, slot, is a approximate value). In the above definition, $w=0.3$, $k_{z}=0.05$, $\epsilon_{z}=12$ and $\epsilon_{q}=1.8$. The center of positional $\mathrm{RBF}$, ${c_{z,i}}$, is evenly chosen on interval $[0,1]$, and the center of orientation $\mathrm{RBF}$, ${\boldsymbol{q}_{c,i}}$, is generated by $(\cos{\frac{\alpha}{2}},\sin{\frac{\alpha}{2}},0,0)*\boldsymbol{q}_{\text{init}}$ with $\alpha$ evenly choosen from $[-\pi/4,3\pi/4]$ to make sure ${\boldsymbol{q}_{c,i}}$ covers all possible orientations. $\beta_{t}=0.9^{t}$ if $t<5$ and 00 otherwise. The bounds of the initial hypothesis space $\Theta_{0}$ is set to $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{h}=[5,\dots,5]^{\mkern-1.5mu\mathsf{T}}$. $\gamma$ is set to 1.5. This initialization leads to a conservative MPC policy, preventing the arm from approaching the target at the beginning.
VI-C2 Human Correction Interface
TABLE II: keyboard commands in robot arm user study
Key Correction Interpretation
"shift" $[0,0,0,-1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go up in world frame
"ctrl" $[0,0,0,1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go down in world frame
"Enter" N/A Reset robot position and switch target.
@@ VI-C2 Human Correction Interface
Figure 13: Illustrations of the drone MPC rollout trajectory after $k=2,6,12$ corrections. The red stars denote the target positions and the blue curves denote the trajectories. The gate is drew in black lines and collision is denoted by green crosses. User’s corrections typically lower the flying drone to make it pass through the gate.
VI-D Participants, Procedures, and Metrics
VI-D1 Participants
We recruited 15 participants from the student population of Arizona State University, covering undergraduate, master’s, and PhD students with different backgrounds. Among all participants, 13 are male and 2 female. All participants are novices to the user study. This user study had been reviewed and approved by the Institutional Review Board (IRB) of Arizona State University, and all participants had signed the consent forms.
VI-D2 Procedure
@@
In the drone navigation game, a trial is deemed successful when the drone navigates through the gate without collision and successfully reaches different targets 6 consecutive times with the learned safe MPC policy. A trial is deemed failed when there are too many corrections, causing the crash of the MVE solver, or the user believes the robot cannot be corrected anymore and gives up.
In the Franka reaching game, a trial is deemed successful when the robot arm reaches different targets without collision with the slot frame with the learned safe MPC controller. The failure of one trial is determined when there are too many corrections or the user believes the robot cannot be corrected anymore and gives up.
The result of every trial is recorded. Before formally starting the study, every user is instructed about the settings of the game and is given 2 chances for hands-on trying of the environments to get familiar with the keys and commands.
VI-D3 Metrics for Evaluation
Two major metrics are used to evaluate the performance of the proposed safe MPC alignment algorithm in our user study:
@@ VI-D3 Metrics for Evaluation
    Numbers of corrections needed for a successful trial: The averaged number of corrections performed by the users in successful trials is a measure for the efficiency of the proposed method. Fewer corrections in successful trials suggest better efficiency in teaching the robot.
VI-E Result and Analysis
VI-E1 User Study Metrics Report
TABLE III: Result of Simulated User Study
@@ VI-E1 User Study Metrics Report
For the user study, we report the success rate and correction count in a successful game trial in Table III. Both metrics are averaged across individual users. The success rate is $89.3\%\pm 7.7\%$ for the drone navigation game and $89.3\%\pm 11.2\%$ for the Franka arm constrained reaching game. The number of corrections for a successful trial is $15.7\pm 10.6$ for the drone navigation game and $16.5\pm 6.1$ Franka arm reaching games.
@@
The above result illustrates that the proposed method is effective (i.e., with a high success rate) in enabling a robot to learn a safety constraint for varying tasks (i.e. reaching targets are changing). It also shows the efficiency of the proposed method — successful learning of a safety constraint only requires tens of human corrections. Since this method is online and the user can perform multiple corrections in one robot rollout trajectory, the constraint can be learned to guide the robot to complete safety-critical tasks in a few runs. In Table III, one can note that the drone navigation task has a higher variance in the correction count than the Franka arm reaching task. This could be attributed to the larger motion range and longer task rollout horizon of drone navigation. As a result, in the drone game, the human correction strategies, including the correction position and correction timing, can vary significantly. Such diverse correction strategies can lead to a larger variance of total correction count.
VI-E2 Analysis for Learned Behavior and Learned Constraint
@@
To demonstrate the learning progress, in Fig. 13 and Fig. 14, we illustratively show the intermediate robot motion at different numbers of human corrections. Fig. 13 shows the flight trajectory of the drone after different numbers of human corrections. Fig. 14 shows the position and altitude of the robot arm’s end effector when passing the slot after the different number of human corrections. From the illustration examples, we can conclude that the learned constraint generally guides the robot to complete the task.
@@
Figure 14: Illustrations of the robot arm pose when passing the $yz$-plane of the bars after $k=4,8,17$ corrections. (a): The motion after 4 corrections with three different targets (cyan ball). (b): The motion after 8 corrections. (c): The motion after 17 corrections. The ends of the gripper are denoted with red balls. At the beginning, the gripper cannot pass the slot. But after several corrections, it succeeds in passing the slot for all target positions.
@@
To better understand the learned safety constraint function ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$, in Fig. 15, we draw the heat map of the value of ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi})$ with $\boldsymbol{\theta}_{H}$ the averaged value of weights learned in all successful trials. In drawing such heat map, since ${g}_{\boldsymbol{\theta}_{H}}$ is defined by RBFs spreading only on $yz$-plane, we fix $r_{x,t}=10$ (recall $x_{\text{gate}}=9.8$) and $r_{y,t},r_{z,t}$ are generated by the grid in $[0,10]\times[0,10]$ for all states $\boldsymbol{x}_{t}$ in $\boldsymbol{\xi}$. In the heat map, we draw the 00-level contour of ${g}_{\boldsymbol{\theta}_{H}}$ in the dashed line, which is the learned safety boundary. Also, we plot the $yz$-plane projection of the gate, which is the black box.
@@
In Fig. 15, we can note the learned safety boundary (corresponding to ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi})\leq 0$) is only similar to the top edge of the gate. This is explained by the following reasons. As shown in Fig. 11, the target positions are set higher than the gate, and thus during the learning process, the motion of the drone is always going up, as shown in Fig. 13. As a result, the majority of human directional corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom boundary untouched. It is expected that in order to learn a safety constraint that fully represents the gate frame, lower target positions should be added so that the drone will receive human corrections near the bottom of the gate.
@@
Figure 15: Heatmap for the learned constraint function in the drone user study. The black solid line denotes the actual position of the gate in $yz$-frame. The black dashed curve denotes the boundary of the safety zone. Since the target positions are higher than the gate (see Fig. 11), thus during the learning process, the motion of the drone is always going up, as shown in Fig. 13. As a result, the majority of human directional corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom boundary untouched.
VII User Study in Real-world Robot Experiment
@@
In this section, we conduct a user study to evaluate the performance of our safe MPC alignment algorithm using a hardware Franka Research 3 robot experiment. We focus on the task of mobile liquid pouring, i.e., the robot performs liquid pouring while moving from one location to another. In certain scenarios, this kind of task can be safety-critical: for example, if the liquid is corrosive/chemical/toxic, inappropriate mobile pouring will cause safety concerns. The task is challenging because it is hard to manually define a safety constraint for the robot to accomplish liquid pouring without spilling. The Franka robot arm will learn such a safe MPC controller from human physical corrections.
VII-A Hardware Experiment Setup
@@
We specifically focus on mobile water pouring by the Franka robot arm. A task setup is shown in Fig. 16. A water pot is gripped by the robot end-effector. The robot needs to move from the starting point $[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ (with orientation quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$) to target point $[-0.65,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$ (with orientation quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$). During the motion, the robot needs to pour more than 50 grams of water into a bar sink. The bar sink is approximately placed in the middle of the robot’s moving path, but the precise size and location of the bar sink are unknown.
Figure 17: Components of the hardware Franka arm learning platform
As shown Fig. 17, the hardware Franka arm platform has three main components: the parameterized safe MPC controller (high-level controller), the low-level controller, and a learning module. The high-level safe MPC controller follows the safe MPC formulation of Franka arm reaching game in Section VI-C (recall that the unknown constraint is parameterized by RBFs). The other control parameters of this Safe MPC controller are listed in Table IV. The high-level safe MPC controller receives the end-effector’s current pose $\boldsymbol{x}$ from the low-level controller and solves for the current desired velocity $\boldsymbol{u}$ (defined in (36)) of the end-effector. The low-level controller implements the operation space velocity control to track the desired velocity $\boldsymbol{u}$ given by the high-level controller.
The learning module updates the parameterize $\boldsymbol{\theta}$ of the high-level safe MPC controller. The update only happens when the learning module detects a human correction $\boldsymbol{a}$, which is always accompanied by an end-effector pose $\boldsymbol{x}_{\boldsymbol{a}}$ where such human correction $\boldsymbol{a}$ is applied to. The learning module implements the proposed safe MPC alignment algorithm.
TABLE IV: parameter list in real-world user study
@@
In the high-level MPC policy, the cost of the MPC controller is set to penalize the current and target gripper position in Cartesian space, as in Section VI-C. Since it is difficult to explicitly define a parameterized constraint using the state of pouring water, we instead define the constraints of safe water pouring using the robot state. Specifically, we define the same RBF (40) parameterized constraint in Section VI-C, constraining the tilt angle and height of the robot gripper for pouring the water into the sink.
Figure 18: Illustrations of human correction in the real-world user study. (a) A human user applies force to the robot. (b) The human moves the robot to the desired configuration to pour the water into the bar sink. The delta position of the end effector is used as human directional correction, as in (42).
As shown in Fig. 18, to give a physical correction to the Franka robot arm, a human user can physically touch/hold any link of the robot arm, and the applied force will immediately activate the zero-torque control mode. Suppose the joint positions of the robot arm before and after human physical correction are $\mathbf{q}_{\text{before}}$ and ${\mathbf{q}_{\text{after}}}$, respectively, we calculate the corresponding raw human correction $\boldsymbol{a}_{r}$ using
@@
with $\mathrm{J}$ the Jacobian of the end-effector. The above calculated $\boldsymbol{a}_{r}\in\mathbb{R}^{6}$ is considered as the directional displacement of the end-effector pose, intended by the human user. In the computed $\boldsymbol{a}_{r}\in\mathbb{R}^{6}$, we mask the entries in the other dimensions with 00s and only keep the entries corresponding to the tilt angle along $x$-axis and the height in $z$-axis, since they are the only dimensions relevant to the given task. The masked vector are is as the actual human correction $\boldsymbol{a}$ fed to the learning module.
@@
Figure 19: Illustrations of one trial of the real-world study. The trajectories of the watering pot is denoted by blue dashed lines. (a) After 2 corrections, the arm begins tilting the pot but spills out the water. (b) After 3 corrections, the arm can pour water into the sink but the human is not satisfied with its height and altitude of pouring. (c) After 5 corrections, the arm learns to tilt only along the direction of the $x-$axis and successfully aligns with human intent.
VII-B Participants, Procedures, and Metrics

### v2

@@ VII-A Setting of Two Simulation Games
Figure 15: Drone navigation game. (a) Tracking view. (b) Fixed view. Users can see the motion of the drone with the help of these two views. The $xyz$-axis (red, green, and blue axes), the gate (yellow bars), the drone, the local plan solved from the current MPC (red curve) and the motion target (yellow balls) are visible to users. In each game reset, the user will see one target randomly.
@@
Figure 16: Franka arm reaching game. (a) The starting pose of the arm. RGB arrows denote the world $xyz$-frame. (b) One example of successful reaching, and long green arrow denotes the motion direction. Users can see the motion of the robot arm by adjusting the view. The slot, the robot arm, and current motion target (cyan ball) are all visible to users.
@@
Figure 17: Illustrations of the drone MPC rollout trajectory after $2,6,12$ corrections. The red stars denote the targets and blue curves denote the trajectories. The gate is drawn as black frame and the collision is highlighted in green crosses. User corrections typically lower the drone to make it pass through the gate.
VII-B Participants, Procedures, and Metrics
VII-B1 Participants
We recruited 15 participants from Arizona State University, including undergraduates, masters, and PhD students with different backgrounds. Among all participants, 13 are male and 2 female. All participants are new to the user study. This user study had been reviewed and approved by the Institutional Review Board (IRB) of Arizona State University, and all participants had signed the consent forms.
VII-B2 Procedure
@@
In the drone navigation game, a trial is deemed successful when the drone navigates through the gate without collision and reaches different targets 6 consecutive times with the learned safe MPC policy. A trial is deemed failed when there are too many corrections (exceeding 25), or the user believes the robot cannot be corrected anymore and gives up.
In the Franka reaching game, a trial is deemed successful when the robot arm reaches different targets with the end-effector having no collision with the slot. The failure of one trial is determined when there are too many corrections (exceeding 25) or the user believes the robot cannot be corrected anymore and gives up.
The result of every trial is recorded. Before formally starting the study, every user is instructed about the settings of the game and is given 2 chances to interact with the environment hands-on to get familiar with the keys.
VII-B3 Metrics for Evaluation
Two major metrics are used to evaluate the performance of the proposed algorithm:
@@ VII-B3 Metrics for Evaluation
    Numbers of corrections needed for a successful trial: The averaged number of corrections performed by users in successful trials is a measure for the efficiency of the proposed method. Fewer corrections in successful trials suggest better human data efficiency.
VII-C Result and Analysis
VII-C1 User Study Metrics Report
TABLE II: Result of user study for simulation games
@@ VII-C1 User Study Metrics Report
For the user study, we report the success rate and correction count in a successful game trial in Table II. Both metrics are averaged across users. The success rate is $89.3\%\pm 7.7\%$ for the drone navigation game and $89.3\%\pm 11.2\%$ for the Franka arm reaching game. The number of corrections for a successful trial is $15.7\pm 10.6$ for the drone navigation game and $16.5\pm 6.1$ Franka arm reaching game. In both games, each learning update of $\boldsymbol{\theta}$ takes approximately 0.15 seconds, during which the simulation briefly pauses. This brief pause is imperceptible to the user, especially since human corrections occur infrequently.
@@
The above result illustrates that the proposed method is effective (i.e., with a high success rate) in learning a safety constraint for varying tasks (recall targets are changing across trials). It also shows the efficiency of the proposed method — successful learning only requires tens of human corrections. Since this method is online and the user can perform multiple corrections in one robot rollout trajectory, the constraint can be learned in a few runs. In Table II, one can note that the drone navigation task has a higher variance in the correction count than the Franka arm reaching task. This could be because the drone game has a larger motion range and longer task rollout horizon. As a result, human correction strategies, such as correction position and timing, can vary significantly. This correction diversity leads to a larger variance of correction count.
VII-C2 Analysis for Learned Constraint and Robot Motion
@@
To illustrate learning progress, in Fig. 17 and Fig. 18, we show the intermediate robot motion at different numbers of human corrections. Fig. 17 shows the trajectory of the drone after different numbers of human corrections. Fig. 18 shows the pose of the robot gripper when passing the slot after a different number of human corrections.
(a) After 4 Corrections
(b) After 17 Corrections
@@
Figure 18: Screenshots of the robot gripper when passing through the slot after $k=4,17$ corrections, given three different targets (cyan ball). (a): After 4 corrections, the gripper collides with the slot. (b): After 17 corrections, the gripper succeeds in passing the slot for all targets.
@@
To better understand the learned safety constraint ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$, in Fig. 19, we draw the heat map of the learned ${g}_{\boldsymbol{\theta}^{*}}(\boldsymbol{\xi})$ for the drone navigation game, with $\boldsymbol{\theta}^{*}$ the averaged of weights learned. Recall in drone navigation game, ${g}_{\boldsymbol{\theta}}$ is defined on y-z plane at $x=10$ (true gate position is $x_{\text{gate}}=9.8$). In the heat map, we draw 00-level contour of ${g}_{\boldsymbol{\theta}^{*}}$ in dashed lines, which is the learned safety boundary. Also, we plot the $yz$-plane projection of the gate, which is the black box.
@@
Fig. 19 shows that the learned safety boundary (corresponding to ${g}_{\boldsymbol{\theta}^{*}}(\boldsymbol{\xi})\leq 0$) is only similar to the top edge of the gate. This is explained by the following reasons. As shown in Fig. 15, the targets are set higher than the gate, and thus during the learning process, the motion of the drone is always going up, as shown in Fig. 17. As a result, the majority of human corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom edge untouched. It is expected that to learn a safety constraint that fully represents the gate, lower target positions should be added so that the drone will receive human corrections near the bottom edge of the gate.
@@
Figure 19: Heatmap for the learned constraint in the drone game. The black solid line denotes the actual position of the gate in $yz$-plane. The black dashed curve denotes the boundary of the learned safety zone. Since the targets are higher than the gate (see Fig. 15), during the learning process, the motion of the drone is always going up, as shown in Fig. 17. As a result, the majority of human directional corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom boundary untouched.
VIII User Study in Real-world Experiment
@@
In this section, we conduct a user study to evaluate the performance of our safe MPC alignment algorithm in a hardware Franka Research 3 experiment. We focus on the task of mobile liquid pouring, i.e., the robot performs liquid pouring while moving from one location to targets. In certain scenarios, this kind of task can be safety-critical: for example, if the liquid is corrosive/chemical/toxic, inappropriate mobile pouring will cause safety concerns. The task is challenging because it is hard to manually define a safety constraint for the robot to accomplish liquid pouring without spilling. The Franka robot arm will learn a safe MPC controller from human physical corrections.
VIII-A Hardware Setup
@@
The setup of mobile water pouring task is shown in Fig. 20. A water pot is gripped by the robot end-effector. The robot moves from the starting point $[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$) to target point $[-0.65,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$). During the motion, the robot needs to pour more than 50 grams of water into a bar sink. The bar sink is placed in the middle of the robot path, but the precise size and location of the bar sink are unknown.
Figure 21: System diagram for the hardware implementation. Different blocks indicate separate processes connected via arrows that indicate communication via ROS.
As shown Fig. 21, the hardware Franka arm platform has three main components: the parameterized safe MPC controller (high-level controller), the low-level controller, and safe MPC alignment module. The safe MPC controller follows the safe MPC formulation of Franka arm reaching game in Appendix B-D (recall that the learnable safety constraint is parameterized by RBFs). The parameters of this Safe MPC controller are listed in Table III. The safe MPC controller receives the end-effector’s current pose $\boldsymbol{x}$ and generate the action $\boldsymbol{u}$ (see (58) in Appendix), which is the desired velocity of the end-effector. The low-level controller implements the operational space control to track the $\boldsymbol{u}$ given by the safe MPC controller.
The safe MPC alignment module updates the parameterize $\boldsymbol{\theta}$ of the safe MPC controller. The update only happens when the alignment module detects a human correction $\boldsymbol{a}$. The module implements our safe MPC alignment algorithm. Note that the alignment module and controller modules runs different process, thus the learning update does not affect the real-time MPC control of the robot arm. ROS is used for communication between different processes.
TABLE III: parameter list in real-world user study
@@
In the safe MPC controller, the cost of the MPC controller is set to penalize the current and target, as in Section B-D. Since it is difficult to explicitly define a parameterized safety constraint using the state of pouring water, we instead define the constraints only using the robot state. Specifically, we use the weighted RBFs to parameterize the safety constraints constraint, similar to Appendix B-D.
As shown in Fig. 20, to give a physical correction to the Franka robot arm, a human user need physically touch/hold any link of the robot arm, and the applied force will immediately activate the zero-torque control mode. Suppose the joint positions of the robot arm before and after human physical correction are $\mathbf{q}_{\text{before}}$ and ${\mathbf{q}_{\text{after}}}$, respectively, we calculate the corresponding raw human correction $\boldsymbol{a}_{r}$ using
@@
with $\mathrm{J}$ the Jacobian of the end-effector. The calculated $\boldsymbol{a}_{r}\in\mathbb{R}^{6}$ is considered as the displacement of the end-effector pose, intended by the human user.
@@
Figure 22: Illustrations of one trial of the real-world experiment. The trajectories of the watering pot is denoted by blue dashed lines. (a) After 2 corrections, the arm begins tilting the pot but spills out the water. (b) After 3 corrections, the arm can pour water into the sink but the human is not satisfied with its height and altitude of pouring. (c) After 5 corrections, the arm learns to tilt only along the direction of the $x-$axis and successfully aligns with human intent.
VIII-B Participants, Procedures, and Metrics